Circular No. 8927

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 E3 (GARRADD) G. J. Garradd reports his discovery of a comet on images taken with the

0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), the object being diffuse and 20′′ in diameter on images from Mar. 7.8 UT. Following posting on the ‘NEOCP’ webpage, E. Guido and G. Sostero (Castellammare di Stabia, Italy) write that they obtained CCD images remotely using a 0.25-m f/3.4 reflector near Moorook, Australia, on Mar. 8.7; thirty co-added 60-s exposures show the comet to have cometary appearance, with a coma diameter of nearly 10′′ with a strong central condensation and a broad tail 18′′ long toward the north.

2008 UT  α_{2000}  δ_{2000}  Mag.
Mar. 5.77805  19^{h}35^{m}19.^{s}03  −41^{o}44′05′′.8  17.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-E70.

T = 2008 Oct. 31.240 TT  ω = 229.^{o}143
Ω = 106.195 }2000.0
q = 5.26127 AU  i = 106.590

NOVA CYGNI 2008 S. Nakano, Sumoto, Japan, reports the discovery by H. Kaneda (Sap-

poro, Japan) of an apparent nova (mag 8.2 ± 0.3) on unfiltered CCD frames taken on Mar. 7.801 UT using a telephoto lens, giving the position of the variable as α = 19^{h}58^{m}33^{s}.4, δ = +29^{o}52′04′′ (equinox 2000.0). Nakano forwards the following precise position end figures for the nova from a CCD exposure on Mar. 8.716 by K. Kadota (Ageo, Japan, 0.25-m reflector): 33^{s}.39, 06′′.5, mag 7.4. Numerous additional apparent independent discoveries were reported on CBET 1291.

D. Nogami and J. Kuriyama, Kyoto University; and I. Iwata, National Astronomical Observatory of Japan, write that a low-resolution (R ∼ 500) spectrum (range 400–800 nm), taken of the new variable on Mar. 8.794 UT with a 1.88-m telescope (+ KOOLS) at the Okayama Astrophysical Observatory, shows a blue continuum with strong Balmer and Fe II lines having prominent P-Cyg profiles. The FWHM of the emission component of Hα is ∼ 1000 km/s, and the minimum of the Hα absorption is blueshifted from the emission peak by ∼ 880 km/s. These features suggest that this object is indeed a classical nova before or around maximum.

2008 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green